Q13.

Suppose T, S : R^2 "\to" R^2 are linear defined by T (u, v) = (3u + v, u + 2v) and S (x, y) = (2x - y, x + y).

Also the matrices of T and S with respect to the standard bases of R^2 and R^2 are given as M (T) ="\begin{bmatrix}
 3 & 1 \\
 1 & 2
\end{bmatrix}"

and M (S) ="\begin{bmatrix}
 2 & - 1 \\
 1 & 1
\end{bmatrix}". Then M(T S) =

(1) "\begin{bmatrix}
 1 & 0 \\
 0 & 1
\end{bmatrix}"

(2) "\begin{bmatrix}
 3 & 1 \\
 1 & 5
\end{bmatrix}"

(3) "\begin{bmatrix}
 3 & 1 \\
 4 & 1
\end{bmatrix}"

4) None of the given answers is true


1
Expert's answer
2022-01-19T18:03:14-0500

M (T S) ="\begin{bmatrix}
 3 & 1 \\
 1 & 2
\end{bmatrix}""\begin{bmatrix}
 2 & - 1 \\
 1 & 1
\end{bmatrix}"="\begin{bmatrix}
 7 & - 2 \\
 4 & 1
\end{bmatrix}"

Answer is (4) None of the given answers is true.